OpenStudy (raffle_snaffle):

Find the angle between the vectors. (First find an exact expression and then approximate to the nearest degree.) a=j+k , b=i+2j-3k

OpenStudy (raffle_snaffle):

I got arccos=(1/2sqrt7) It's suppose to be a negative one instead though... I can't find the problem unless the book is wrong.

OpenStudy (raffle_snaffle):

wait a min I think I know what I did wrong...

OpenStudy (raffle_snaffle):

Is i=0?

terenzreignz (terenzreignz):

What do you mean? for the first vector, \(\large\vec a\) the i-component is zero...

OpenStudy (raffle_snaffle):

Okay then I know what I did wrong... I am dyslexic. Lol

terenzreignz (terenzreignz):

Well... just to get you started, the cosine of the angle between two vectors is given by \[\Large \cos(\theta) = \frac{\vec u \cdot \vec v}{||\vec u || \ ||\vec v ||}\]

OpenStudy (amistre64):

i tend to stack, then multiply down the cols, and add the results for a dot product a=0+j+k , b=i+2j-3k ----------- 2 - 3 = -1 right?

terenzreignz (terenzreignz):

Whatever works :D

terenzreignz (terenzreignz):

\[\Large ||\vec a || = \sqrt 2\]\[\Large \left|\left|\vec b\right|\right|=\sqrt {14}\]

OpenStudy (amistre64):

i thought is was |b| = sqrt(2*7) :)

terenzreignz (terenzreignz):

Oh, silly me, how could I have missed that? D: <slashes wrists>

OpenStudy (raffle_snaffle):

Okay thanks guys. I was looking at j as i and made j=0 so that must have screwed me up somewhere...

terenzreignz (terenzreignz):

It'd be nice if you show your answer...

OpenStudy (raffle_snaffle):

Magnitude of a=sqrt2 and magnitude b=sqrt14

terenzreignz (terenzreignz):

No, I mean your final answer as to the angle between the two vectors.

OpenStudy (raffle_snaffle):

ohh Haha

OpenStudy (raffle_snaffle):

theta=arc cos(-1/2sqrt17) =100.893* which is about 101*

terenzreignz (terenzreignz):

...trying to visualise in my head...

terenzreignz (terenzreignz):

That does seem about right :D

OpenStudy (raffle_snaffle):

since a*b=-1 the angle is going to be greater than 90*

terenzreignz (terenzreignz):

Okay, it's correct. Well done ^_^

OpenStudy (raffle_snaffle):

If it was 1 then the angle would be less than 90*

OpenStudy (amistre64):

cos is negative in q2 in q3 yes, and positive in q1 and q4

OpenStudy (amistre64):

cos(90) = 0

terenzreignz (terenzreignz):

I don't know these shortcuts.. D:

OpenStudy (amistre64):

well if youd spend less time slashing, and more time nose grinding ...

OpenStudy (amistre64):